# How to set AdamW’s weight decay as you scale model and dataset size

### v1

@@ A.3.1 Output-batchnorm for scale-invariant output weights
As such, we applied a “global batchnorm” layer after the output layer, i.e., upon the logits. In particular, denote the logits by $Z\in\mathbb{R}^{B\times C}$, where $B$ is the batchsize and $C$ is the number of classes, we first flatten (and reshape) $Z$ into a matrix of shape $BC\times 1$ and then feed the vector into a standard 1D batchnorm, which computes the mean and variance across all $BC$ elements. Finally, we reshape the output of batchnorm back to $B\times C$ and send it to the final softmax layer. The effect of output-batchnorm is shown in row C in Fig. 6 and 7.
@@ A.3.3 Decoupling the learning rates for the batch/layernorm parameters
It turns out that we already do something different with these parameters. In particular, it is common practice [^3] We cannot find literature explicitly studying the effect of dropping weight decay for normalization layers’ parameters, however, its importance has been noticed in vision tasks, e.g. by Jia et al. (2018) and https://github.com/JiahuiYu/slimmable_networks/issues/15, as well as in language model training: https://github.com/karpathy/minGPT/blob/master/mingpt/model.py#L227 to optimize them using the same learning rate, $\eta$, as the other parameters, but to drop weight decay. This standard setting is depicted in Fig. 6E and Fig. 7E, and we can see that $\eta$ does still change the behavior of the optimizer.
@@ Appendix B The weights in an EMA
Appendix C Additional experiment results
@@
Appendix D Note on muP library
We used the mup library [^4] https://github.com/microsoft/mup from the authors of Yang et al. (2022) for varying-model-size experiments. In particular, for ResNet experiments in Fig. 3 and 4, we directly use the ResNet codebase under the examples folder together with the provided MuAdamW optimizer to run our experiments, as this library takes care of details such as the scaling of the learning rate and initialization for the input/output layers. For ViT experiments in Fig. 10, we manually constructed the required model shape file using the provided make_base_shapes function and used the provided MuReadout module as the classification head.
@@ Appendix D Note on muP library
Appendix E Extended experiment setups
E.1 Model specification
For ResNet-18 experiments, we utilized the implementation from https://github.com/kuangliu/pytorch-cifar/. For both CIFAR-10 and ImageNet, we used random cropping and horizontal flip as data augmentation and we used cross-entropy as the loss function.
@@ E.1 Model specification
E.2 Hyperparameter range
In Sec. 4.1, for each given $\eta$, we considered, $\eta\lambda$ of
$\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda\in\{10^{-7},3\times 10^{-7},10^{-6},3\times 10^{-6},\ldots,3\times 10^{-2},10^{-1}\}.$ (48)
@@ E.2 Hyperparameter range
Appendix F Licenses

### v2

@@ A.3.1 Output-batchnorm for scale-invariant output weights
Therefore, we applied a “global batchnorm” layer after the output layer, i.e., upon the logits. In particular, denote the logits by $Z\in\mathbb{R}^{B\times C}$, where $B$ is the batchsize and $C$ is the number of classes, we first flatten (and reshape) $Z$ into a matrix of shape $BC\times 1$ and then feed the vector into a standard 1D batchnorm, which computes the mean and variance across all $BC$ elements. Finally, we reshape the output of batchnorm back to $B\times C$ and send it to the final softmax layer. The effect of output-batchnorm is shown in row C in Fig. 7 and 8.
@@ A.3.3 Decoupling the learning rates for the batch/layernorm parameters
It turns out that we already do something different with these parameters. In particular, it is common practice333 We cannot find literature explicitly studying the effect of dropping weight decay for normalization layers’ parameters, however, its importance has been noticed in vision tasks, e.g. by Jia et al. (2018) and https://github.com/JiahuiYu/slimmable_networks/issues/15, as well as in language model training: https://github.com/karpathy/minGPT/blob/master/mingpt/model.py#L227 to optimize them using the same learning rate, $\eta$, as the other parameters, but to drop weight decay. This standard setting is depicted in Fig. 7E and Fig. 8E, and we can see that $\eta$ does still change the behavior of the optimizer.
@@ Appendix B The weights in an EMA
Appendix C Weight norm
Notice that the EMA perspective of AdamW (Eq. (6)) suggests that AdamW averages over recent $q_{t}=-\frac{1}{\lambda}\tfrac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$, and if we assume $\tfrac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ is of scale $\mathcal{O}(1)$ (Balles & Hennig, 2018), the weights acquired from AdamW should be of magnitude $\mathcal{O}(\frac{1}{\lambda})$.
To verify this, we trained a ResNet on CIFAR-10 for 1000 epochs under a batch size of $100$, with an initial timescale $\tau_{\text{iter}}=1/(\eta_{0}\lambda)=10^{5}$. We used a cosine learning rate schedule, starting from $\eta_{0}=10^{-5}/\lambda$ and decaying to 00. We then swiped $\lambda$ between $\{10^{-3},3\times 10^{-3},10^{-2},\ldots,0.3,1.0\}$ and recorded the average absolute value of all dimensions in the convolutional and linear layer weights (i.e. weights where we adopt weight decay) under different $\lambda$.
Fig. 11 left plots the final weight magnitudes against $\lambda$, with both X and Y axes log-scaled. By comparing it with the reference line $y=0.01/x$ (gray, dashed), we can see the magnitude against $\lambda$ line shares the same slope, indicating that the weight magnitude approximately follows $\mathbb{E}[|W_{i}|]=c/\lambda$, where $c$ is a constant. Fig. 11 right plots the weight magnitudes throughout optimization, where we can see that all configurations start with the same magnitude at initializaiton, but gradually converge to our predicted magnitudes as optimization proceeds.
Figure 11: Weight magnitudes scale inversely with $\lambda$ AdamW optimization. We trained a ResNet-18 on CIFAR-10 using AdamW with cosine learning rate decay over 1,000 epochs. The weight decay parameter $\lambda$ varies from $10^{-3}$ to $1$, with initial learning rate $\eta=10^{-5}/\lambda$. Left: Final average absolute weight values $\mathbb{E}[|W_{i}|]$ in convolutional and linear layers versus $\lambda$, showing approximate $1/\lambda$ scaling (same slope as the reference dashed gray line). Right: Evolution of weight magnitudes throughout training for different $\lambda$ values, different runs start with the same magnitude and gradually converge to $\mathcal{O}(1/\lambda)$.
Appendix D Additional experiment results
D.1 ViT scaling experiments on ImageNet
@@
Figure 12: Optimal $\lambda$ and $\tau_{\text{epoch}}$ v.s. dataset sizes under a constant learning rate. Similar to Fig. 1 but use a constant learning rate of $10^{-3}$ rather than cosine decay to 0.1 of the initial learning rate.
Figure 13: Optimal $\lambda$ and $\tau_{\text{epoch}}$ v.s. dataset sizes under a cosine decay schedule to 0. Similar to Fig. 1 but use a cosine learning rate decay schedule from $10^{-3}$ to $o$ rather than to $10^{-4}$.
@@
Figure 15: Similar setting to Fig. 3, but use constant learning rate.
Figure 16: Similar setting to Fig. 3, but use cosine decay to zero.
Figure 17: Similar setting to Fig. 5, but use constant learning rate.
Figure 18: Similar setting to Fig. 5, but use cosine decay to zero.
Appendix E Note on muP library
We used the mup library [^4] https://github.com/microsoft/mup from the authors of Yang et al. (2022) for varying-model-size experiments. In particular, for ResNet experiments in Fig. 3 and 5, we directly use the ResNet codebase under the examples folder together with the provided MuAdamW optimizer to run our experiments, as this library takes care of details such as the scaling of the learning rate and initialization for the input/output layers. For ViT experiments in Fig. 14, we manually constructed the required model shape file using the provided make_base_shapes function and used the provided MuReadout module as the classification head.
@@ Appendix E Note on muP library
Appendix F Extended experiment setups
F.1 Model specification
For ResNet-18 experiments, we utilized the implementation from https://github.com/kuangliu/pytorch-cifar/. For both CIFAR-10 and ImageNet, we used random cropping and horizontal flip as data augmentation and we used cross-entropy as the loss function. For µP experiments with ResNet, we use the codebase  [^6] https://github.com/microsoft/mup/tree/main/examples/ResNet provided by Yang et al. (2022), which is an adaptation of the codebase provided by the GitHub user kuangliu.
@@ F.1 Model specification
For dataset size scaling experiments with NanoGPT in Fig. 2, we used the configuration code provided by  (D’Angelo et al., 2024) [^7] https://github.com/tml-epfl/why-weight-decay/blob/main/large_language_models/config/train_gpt2_small_block256.py. The model structure is identical to the standard 124M version of NanoGPT, which contains 12 layers, each layer has 12 attention heads with a head dimension of 64, i.e. an embedding size of 768. Identical to the original NanoGPT, (D’Angelo et al., 2024) uses a micro-batch size of 12 sequences and 40 gradient accumulation steps at each iteration. However, to reduce memory consumption, it uses a fixed context length of 256 tokens instead of the original 1024-token context length.
For width transfer experiments in Fig. 4, we used the codebase from the recent project nanoGPT-mup [^8] https://github.com/EleutherAI/nanoGPT-mup. In particular, we trained transformers with 8 layers, with a fixed attention head dimension of 64, and varied the number of attention heads between $\{4,8,16\}$. We additional included QK layernorm for better stability when swiping hyperparameters. The model is trained with a micro-batch size of 8 sequences, 16 gradient accumulation steps, and a fixed context length of 256 tokens.
F.2 Hyperparameter range
@@
For the NanoGPT experiments in Fig. 2, we used $\lambda=2^{i}$ with $i\in\{-8,-7,\ldots,1\}$.
For the NanoGPT experiments in Fig. 5. We used $\lambda_{\text{base}}=2^{i}$ with $i\in\{-8,-7,\ldots,-1\}$ for $\eta_{\text{base}}=3\times 10^{-4}$ and $\lambda_{\text{base}}=2^{i}$ with $i\in\{-6,-5,\ldots,1\}$ for $\eta_{\text{base}}=3\times 10^{-5}$
Appendix G Licenses